OpenStudy (anonymous):

How can I Integrate (secx+tanx)^2?

OpenStudy (shubhamsrg):

=> sec^2 x + tan^2 x +2 tanx sec x now you can integrate each term separately integral(sec^2 x)= tanx write tan^2 x = sec^2 x -1 and integral(sec x tanx) = sec x..

OpenStudy (anonymous):

According to Wolfram, you get \[ \int (\tan (x)+\sec (x))^2 \, dx=\frac{4 \sin \left(\frac{x}{2}\right)}{\ \cos \left(\frac{x}{2}\right)-\sin \left(\frac{x}{2}\right)}-x + C \]

OpenStudy (anonymous):

Ah! I got it! Thanks guys!